OpenStudy (anonymous):

Help please! Use the Fundamental Theorem of Calculus to find the derivative of... h(x)= integral of -3 to sin(x) (cos(t^3)+sin(t))

OpenStudy (anonymous):

I know i have to use the rule h(x)= integral of a to u(x) g(t) dt equals h'(x)= u'(x)g(u(x))

OpenStudy (anonymous):

let f(x)=cos(t^3)+sin(t) then h(x)=F(sin(x))-F(-3) then h'(x)=d/dx(F(sin(x)))-0=F'(sin(x))*sin'(x)=f(sin(x))*cos(x)

OpenStudy (anonymous):

(F is the antiderivative of f)

OpenStudy (anonymous):

right so that would leave the equation to be... (cos(sin^3(x))+sin(x))+(cos(-3)+(-3))

OpenStudy (anonymous):

f(sin(x))=cos(sin^3(x))+sin(sin(x)) f(sin(x))*cos(x)=cos(sin^3(x))*cos(x)+sin(sin(x))*cos(x)

OpenStudy (anonymous):

i am so confused :(

OpenStudy (anonymous):

using the rule h(x) = integral of a to u(x) g(t) dt equals h'(x)= u'(x)g(u(x)) a=-3 u(x)=sin(x) g(t)=cos(t^3)+sin(t) now find h(x)

OpenStudy (anonymous):

*now find h'(x)

OpenStudy (anonymous):

h'(x)= u'(x)g(u(x)) u'(x)=cos(x) g(u(x))=g(sin(x))=cos(sin(x)^3)+sin(sin(x)) h'(x)=u'(x)g(u(x))=(cos(sin(x)^3)+sin(sin(x))) * cos(x)

OpenStudy (anonymous):

where did the extra sin for g(u(x)) come from?

OpenStudy (anonymous):

I just have no idea what to do :(